# TerraCodec: Compressing Earth Observations

### v1

@@ 4.2 Evaluation
Baselines. We compare TerraCodec to widely used classical codecs: JPEG (Wallace, 1991), JPEG2000 (Taubman & Marcellin, 2002), WebP (Zern et al., 2024), and HEVC (x265) (Sullivan et al., 2012), using the highest bit support for each codec. For JPEG and WebP we use the Pillow library; for JPEG2000 we use Glymur; and for HEVC we use the ffmpeg x265 implementation with the medium preset tuned for PSNR. Since these codecs are limited to three-channel RGB, we apply them per band, encoding each spectral channel as an independent grayscale image, following prior work in EO compression (Radosavljevic et al., 2020; Grassa et al., 2022).
Metrics. Compression rate is reported as bits-per-pixel-band-frame (bppbf), which normalizes by spatial resolution, number of spectral bands, and sequence length. Unlike the standard bits-per-pixel (bpp) used for three-channel RGB images, bppbf extends to arbitrary channel counts and sequence lengths. Distortion is quantified using PSNR, SSIM (Wang et al., 2004), MS-SSIM (Wang et al., 2003), and MSE in the destandardized 16-bit reflectance space. To ensure equal contribution of all spectral channels, metrics are computed per band and averaged across channels, avoiding bias toward bands with higher variance.
@@ 4.3 Downstream tasks
Cloud inpainting. TEC-TT captures spatiotemporal priors that can be applied beyond compression. We demonstrate this with zero-shot cloud removal on the AllClear benchmark (Zhou et al., 2024). Each sample consists of three cloudy observations with masks and one cloud-free target. To apply TEC-TT, we use the two least cloudy images as a temporal context and extract all cloud-free patches from the least cloudy image as input $x_{0}$. Patches in $x_{0}$ that are covered by clouds are predicted by TEC-TT. We report PSNR and other metrics, comparing against the official AllClear baselines. Following the benchmark, metrics are computed in auto mode across all bands, not per-band.
Downstream models on compressed data. To assess the impact of compression on subsequent processing, we finetune AI models on reconstructed inputs from TEC-FP. We evaluate on Sen1Floods11 (Bonafilia et al., 2020), consisting of $512^{2}$ patches from 11 flood events, with binary segmentation masks. The original dataset includes Sentinel–2 L1C imagery. We, therefore, redownload the L2A version for TEC-FP. For patchwise multi-label land cover classification, we use reBEN-7k (Marti-Escofet et al., 2025), which spans eight countries and 19 semantic labels. Images have $120^{2}$ pixels; we apply reflect padding to match the model input sizes. We compress all inputs at three different operating points using TEC-FP, then fine-tune pretrained models on the reconstructed data. We employ Prithvi 2.0 100M (Szwarcman et al., 2025) and TerraMind base (Jakubik et al., 2025) backbones, with a UNet decoder for segmentation and a linear head for classification. All models are trained for 100 epochs with AdamW ( $\text{lr}\!=\!5*10^{-5}$) and a reduce-on-plateau scheduler.
@@ 5 Experiments
We evaluate TerraCodec in terms of rate–distortion (Section 5.1), flexible-rate compression (Section 5.2), and utility for downstream EO tasks (Section 5.3).
@@ 5.1 Rate–distortion
Figure 5: Rate–distortion curves for PSNR (↑) and SSIM (↑) on SSL4EO-S12 v1.1 validation sequences. TerraCodec models consistently outperform standard codecs, with TEC-TT achieving the best overall performance by exploiting temporal context.
@@
This limits temporal gains since half of the frames are bootstrap frames, which lack full previous context, compared to P-frames, which are predicted from two past frames. To better understand these effects, we study the role of temporal conditioning and report the RD performance on P-frames only in Figure 6. Without context, TEC-TT reduces to an image codec (labeled as image only). For a medium setting ($\lambda=5$), conditioning on one past frame improves compression by 13.6% compared to no context. On P-frames with two previous images, TEC-TT achieves a 22.6% rate reduction at equal PSNR, showing that longer EO sequences naturally yield greater efficiency as bootstrap frames are amortized.
@@ 5.2 Flexible rate–distortion
Our TEC-TT-based FlexTEC model uses Latent Repacking to provide flexible rate-compression from a single checkpoint by transmitting a variable subset of latents and inferring the remainder from the model prior. Figure 8 compares FlexTEC against our fixed-rate models and standard codecs on P-frame compression. While fixed-rate TEC-TT models serve as an upper bound—each being optimally fitted for one specific rate–distortion—they require several separately trained models. In contrast, FlexTEC provides several user-controlled rate settings and performs close to or better than TEC-FP, depending on the setting. Further analysis in Appendix E.3 shows that FlexTEC encodes significant information in bootstrap frames, leading to similar bitrates independent of the token budget. The model then uses this information in the following P-frames to provide efficient rate–distortion settings. Figure 8 provides qualitative examples for different token budgets.
@@ 5.3 Downstream tasks
Model beliefs.
Neural codecs rely on learned priors to estimate latent distributions for entropy coding. By decoding these priors into the image space, we obtain predictions that expose the model’s implicit knowledge. Figure 9 shows qualitative examples, with additional results in Appendix F.2.2. We compare priors under three conditions: (a) no information from the current frame (only past context, TEC-TT), (b) partial information from the current latent (past frames and a subset of tokens for TEC-TT; subsets of channel groups for TEC-ELIC), and (c) full information of the latent. TEC-FP, lacking a hyperprior, cannot adapt to the latent at hand. Results show that TEC-TT already produces plausible forecasts from past frames alone and improves its beliefs when partial information becomes available, yielding the most refined predictions.
@@
Cloud inpainting.
Building on TEC-TT’s latent predictions, we evaluate zero-shot cloud removal on the AllClear benchmark (Zhou et al., 2024). TEC-TT is applied without task-specific training (see Sec. 4.3 and Appendix G), with results summarized in Table 1 and qualitative examples in Figure 10. In addition to the full test set, we report performance on subsets ranked by cloud coverage.
The benchmark compares against heuristic baselines (LeastCloudy, Mosaicing) and prior zero-shot neural approaches. TEC-TT outperforms all heuristic methods and prior zero-shot neural models on the full test set. The subset analysis, focusing on the most challenging samples in terms of cloud coverage, further underscores the benefit of temporal priors: heuristic approaches perform adequately in low-cloud cases, but break down under heavy cloud coverage, while TEC-TT maintains high PSNR. The hardest 10% of samples correspond to an average cloud coverage of 99%, yet TEC-TT still produces reasonable predictions. Overall, the results demonstrate that the temporal modeling in TEC-TT not only improves compression but also transfers to challenging forecasting tasks.
Table 1: Test PSNR on AllClear across difficulty subsets (by average cloudiness). PSNR computed across all bands following AllClear.
Model 10% 20% 50% 100%
@@
CTGAN 25.58 26.60 27.59 27.79
DiffCR 24.55 25.13 25.50 25.21
PMAA 24.82 25.06 25.02 24.32
U-TILISE 13.20 14.95 18.33 24.67
UnCRtainTS 26.42 26.50 27.97 29.01
TerraCodec-$T$ 25.97 26.59 30.38 32.86
Figure 10: Cloud inpainting examples with TEC-TT on the AllClear benchmark.
Downstream models.
@@
Table 2: Test performance on reBEN-7k and Sen1Floods11 (↑) when training on compressed inputs (TEC-FP). Numbers in parentheses show the change relative to training on uncompressed data. Performance remains stable at low and mid rates, with clearer degradation for high compression.
@@
Task Model Compression Accuracy F1 mIoU IoUFlood
TerraMind base Original data 88.76 61.99 87.77 78.75
@@ 6 Conclusion
We introduce and release TerraCodec, a family of learned compression models for Earth observation, pretrained on Sentinel-2. Our models outperform classical image and video codecs in rate–distortion, achieving up to an order-of-magnitude reduction at equal quality. Latent Repacking further enables flexible-rate transformer models from a single checkpoint, as demonstrated by FlexTEC. Downstream evaluations show that moderate compression preserves analysis performance, while zero-shot cloud inpainting highlights the strengths of our temporal transformer TEC-TT beyond compression. Overall, TerraCodec establishes a foundation for high-performance multispectral and temporal learned compression in Earth observation.
@@ Acknowledgments
Appendix A Appendix
A.1 Limitations
While TerraCodec is pretrained on a large and diverse global dataset, the current models are limited to Sentinel-2 L2A imagery. Extending the framework to other sensors or future processing versions would require retraining, which we regard as a promising direction for developing cross-sensor codecs. Our implementation, which will be open-sourced, prioritizes research clarity over optimized inference speed; for instance, efficiency techniques such as KV caching are not yet integrated. Incorporating such improvements could further enhance the models’ practicality for real-world deployment. Finally, our flexible-rate variant, FlexTEC, establishes a strong baseline for variable-rate compression using a single checkpoint. While it already performs competitively across many settings, there remains room to further narrow the gap to fixed-rate models, particularly under very high compression ratios.
A.2 Use of LLMs
Appendix B Technical implementation details
@@
B.1 Framework and environment
All models are implemented in PyTorch, using CompressAI (Bégaint et al., 2020) for core architectures and entropy bottlenecks, extended to support multispectral inputs, temporal samples, and model-belief analyses. The TEC-TT implementations are based on Muckley et al. (2021). Experiments are run on NVIDIA A100 GPUs with mixed precision training.
B.2 Image models (TEC-FP, TEC-ELIC)
TerraCodec-FP follows the factorized prior baseline (Ballé et al., 2017), with $g_{a}$ and $g_{s}$ implemented as four strided $5{\times}5$ convolutions combined with GDN/IGDN nonlinearities.
TerraCodec-ELIC builds on the uneven channel-group entropy model (He et al., 2022), using the Chandelier reimplementation (Chandelier, 2023). Latents are divided into groups [16,16,32,64,M128][16,16,32,64,M{-}128] and coded sequentially using SCCTX (space–channel context). Relative to He et al. (2022), we omit the preview head for efficiency. Despite the stronger context model, training time remains comparable to TEC-FP since the channel-group autoregression is parallelizable across spatial locations and implemented with efficient masked convolutions.
@@ B.2 Image models (TEC-FP, TEC-ELIC)
Temporal sampling.
For image models, we treat the pretraining data as an image dataset by sampling individual timesteps from the SSL4EO-S12 time series. During training, a single temporal index is randomly drawn for every sample in each epoch, such that one epoch covers one quarter of the temporal data. Over the 100 training epochs, this amounts to about 25 full passes through the complete dataset.
B.3 Temporal model (TEC-TT)
We provide additional details on the TEC-TT architecture, including its tokenization strategy and temporal transformer design, following VCT (Mentzer et al., 2022).
The image encoder–decoder follows the TEC-ELIC backbone, composed of residual bottleneck and attention blocks with total downsampling $\times 16$. For our $256{\times}256$ input crops, this yields a $16{\times}16$ latent grid with M=192M{=}192 channels.
Tokenization.
We tokenize the latent image representations using the scheme introduced in VCT. The latent grid $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{H_{\ell}\times W_{\ell}\times d_{\text{lat}}}$ with $H_{\ell}=W_{\ell}=16$ is divided into spatial blocks. The current frame is split into non-overlapping $4{\times}4$ blocks, each flattened into a sequence of T=16T{=}16 tokens $\{\hat{\mathbf{y}}_{i,b,t}\}_{t=1}^{16}$. The two past frames are partitioned into overlapping $8{\times}8$ context blocks using reflect-padding so their grids align with the current frame, producing T=64T{=}64 tokens per block.
Concretely, for block index $b\in\{1,\dots,B\}$:
@@ B.3 Temporal model (TEC-TT)
Temporal transformer stack.
The temporal model follows the VCT design and consists of two separate encoders $E_{\text{sep}}$ (one per past frame), a joint encoder $E_{\text{joint}}$ that fuses both contexts, and a masked decoder that autoregressively models the current block tokens conditioned on the fused context. We adopt the standard VCT specifications for the number of layers, heads, and embedding size in each transformer (see Table 4).
@@
Training.
Training follows the standard uniform-noise quantization surrogate (Toderici et al., 2016; Minnen et al., 2018), while inference applies hard quantization and arithmetic coding. We train six TEC-TT variants at different rate–distortion trade-offs, controlled by $\lambda\in\{0.4,5.0,20.0,100.0,300.0,700.0\}$, where smaller values enforce higher compression and larger values prioritize reconstruction quality.
Appendix C Latent Repacking and FlexTEC

### v2

@@ 4.2 Evaluation
Baselines. We compare TerraCodec to widely used classical codecs: JPEG [55], JPEG2000 [50], WebP [62], and HEVC (x265) [48], using the highest bit support for each codec. For JPEG and WebP we use the Pillow library; for JPEG2000 we use Glymur; and for HEVC we use the ffmpeg x265 implementation with the medium preset tuned for PSNR. Since these codecs are limited to three-channel RGB, we apply them per band, encoding each spectral channel as an independent grayscale image, following prior work in EO compression [43, 24].
Metrics. Compression rate is reported as bits-per-pixel-band-frame (bppbf), which normalizes by spatial resolution, number of spectral bands, and sequence length. Unlike the standard bits-per-pixel (bpp) used for three-channel RGB images, bppbf extends to arbitrary channel counts and sequence lengths. Distortion is quantified using PSNR, SSIM, MS-SSIM, and MSE in the destandardized 16-bit reflectance space. To ensure equal contribution of all spectral channels, metrics are computed per band and averaged across channels, avoiding bias toward bands with higher variance.
@@ 4.3 Downstream tasks
Cloud inpainting. TEC-TT captures spatiotemporal priors that can be applied beyond compression. We demonstrate this with zero-shot cloud removal on the AllClear benchmark [66]. Each sample consists of three cloudy observations with masks and one cloud-free target. To apply TEC-TT, we use the two least cloudy images as a temporal context and extract all cloud-free patches from the least cloudy image as input $x_{0}$. Patches in $x_{0}$ that are covered by clouds are predicted by TEC-TT. We report PSNR and other metrics, comparing against the official AllClear baselines. Following the benchmark, metrics are computed in auto mode across all bands, not per-band.
Downstream models on compressed data. To evaluate the effect of compression on downstream analysis, we benchmark downstream task models on compressed–reconstructed versus uncompressed inputs. Following standard EO practice, we finetune task-specific encoder–decoder models on either reconstructed or original inputs. We evaluate on Sen1Floods11 [10], consisting of $\times512$ patches from 11 flood events, with binary segmentation masks. The original dataset includes Sentinel–2 L1C imagery. We, therefore, redownload the L2A version for TEC-FP. For patchwise multi-label land cover classification, we use reBEN-7k [38], which spans eight countries and 19 semantic labels. Images have $\times120$ pixels; we apply reflect padding to match the model input sizes. We compress all inputs using TEC-FP at three different operating points, then fine-tune pretrained models on the reconstructed data. We employ Prithvi 2.0 100M [49] and TerraMind base [29] backbones, with a UNet decoder for segmentation and a linear head for classification. All models are trained for 100 epochs with AdamW ($\text{lr}\!=\!5*10^{-5}$) and a reduce-on-plateau scheduler.
@@ 5 Experiments
We evaluate TerraCodec in terms of rate–distortion (Sec. 5.1), flexible-rate compression (Sec. 5.2), and utility for downstream EO tasks (Sec. 5.3).
@@ 5.1 Rate–distortion
Figure 5: Rate–distortion curves for PSNR (↑) and SSIM (↑) on SSL4EO-S12 v1.1 validation sequences. TerraCodec models outperform standard codecs, with TEC-$T$ achieving best overall performance by exploiting temporal context.
JPEG2000 achieves competitive PSNR with only a 3x lower compression rate at similar distortion—surpassing HEVC—but performs poorly in SSIM, especially at high quality. This arises from its tendency to preserve pixel averages (favored by PSNR) while smoothing textures and edges that SSIM is sensitive to. In contrast, TerraCodec maintains a strong performance across both metrics, demonstrating efficient compression without loss of high-frequency details.
@@
This limits temporal gains because half of the frames are bootstrap frames, which lack full previous context, whereas P-frames are predicted from two past frames. To better understand these effects, we study the role of temporal conditioning and report the RD performance on P-frames only in Figure 6. Without context, TEC-TT reduces to an image codec (labeled as image only). For a medium setting ($\lambda=5$), conditioning on one past frame improves compression by 13.6% compared to no context. On P-frames with two previous images, TEC-TT achieves a 22.6% rate reduction at equal PSNR, showing that longer EO sequences yield greater efficiency as bootstrap frames are amortized.
@@ 5.2 Flexible rate–distortion
Our TEC-TT-based FlexTEC model uses Latent Repacking to provide flexible rate-compression from a single checkpoint by transmitting a variable subset of latents and inferring the remainder from the model prior. Figure 8 compares FlexTEC against our fixed-rate models and standard codecs on P-frame compression. While fixed-rate TEC-TT models serve as an upper bound—each being optimally fitted for one specific rate–distortion—they require several separately trained models. In contrast, FlexTEC provides several user-controlled rate settings and performs close to or better than TEC-FP, depending on the setting. Further analysis in the supplementary material shows that FlexTEC encodes significant information in bootstrap frames, leading to similar bitrates independent of the token budget. The model then uses this information in the following P-frames to provide efficient rate–distortion settings. Figure 8 provides qualitative examples for different token budgets.
@@ 5.3 Downstream tasks
Model beliefs. Neural codecs rely on learned priors to estimate latent distributions for entropy coding. By decoding these priors into the image space, we obtain predictions that expose the model’s implicit knowledge. Figure 9 shows qualitative examples, with additional results in the supplementary material. We compare priors under three conditions: (a) no information from the current frame (only past context, TEC-TT), (b) partial information from the current latent (past frames and a subset of tokens for TEC-TT; subsets of channel groups for TEC-ELIC), and (c) full information of the latent. TEC-FP, lacking a hyperprior, cannot adapt to the latent at hand. Results show that TEC-TT already produces plausible forecasts from past frames alone and improves its beliefs when partial information becomes available, yielding the most refined predictions.
@@
Cloud inpainting. Building on TEC-TT’s latent predictions, we evaluate zero-shot cloud removal on the AllClear benchmark [66]. TEC-TT is applied without task-specific training (see Sec. 4.3 and the supplementary material), with results summarized in Table 1 and qualitative examples in Figure 10. In addition to the full test set, we report performance on subsets ranked by cloud coverage. The benchmark compares against heuristic baselines (LeastCloudy, Mosaicing) and prior zero-shot neural approaches. TEC-TT outperforms all heuristic methods and prior zero-shot neural models on the full test set.
Table 1: Test PSNR (↑) and SSIM (↑) on AllClear across difficulty subsets (by average cloudiness). Computed over all bands following AllClear.
10% 20% 50% 100% (all)
Model PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
@@
CTGAN 25.58 0.765 26.60 0.794 27.59 0.056 27.79 0.840
DiffCR 24.55 0.716 25.13 0.739 25.50 0.758 25.21 0.744
PMAA 24.82 0.746 25.06 0.758 25.02 0.770 24.32 0.768
U-TILISE 13.20 0.546 14.95 0.598 18.33 0.693 24.67 0.807
UnCRtainTS 26.42 0.813 26.50 0.826 27.97 0.057 29.01 0.898
TerraCodec-$T$ 25.97 0.814 26.59 0.830 30.38 0.887 32.86 0.917
Figure 10: TEC-$T$ cloud inpainting examples from AllClear.
The subset analysis, focusing on the most challenging samples in terms of cloud coverage, further underscores the benefit of temporal priors: heuristic approaches perform adequately in low-cloud cases, but break down under heavy cloud coverage, while TEC-TT maintains high PSNR and SSIM. The hardest 10% of samples correspond to an average cloud coverage of 99%, yet TEC-TT still produces reasonable predictions. Overall, the results demonstrate that the temporal modeling in TEC-TT not only improves compression but also transfers to challenging forecasting tasks.
@@
Table 2: Test performance on reBEN-7k and Sen1Floods11 when training on compressed inputs (TEC-FP). Numbers in parentheses show the change relative to training on uncompressed data. Performance remains stable at low and mid rates, with more pronounced degradation for high compression.
@@
Task Model Compression Acc. F1 mIoU IoUFlood
TerraMind 1.0 base Original data 88.76 61.99 87.77 78.75
@@ 6 Conclusion
We introduce and release TerraCodec, a family of learned compression models for Earth observation, pretrained on Sentinel-2 multispectral time series data. Our models outperform classical image and video codecs in rate–distortion, achieving up to an order-of-magnitude reduction at equal quality. Latent Repacking further enables flexible-rate transformer models from a single checkpoint, as demonstrated by FlexTEC. Downstream evaluations show that moderate compression preserves analysis performance, while zero-shot cloud inpainting highlights the strengths of our temporal transformer TEC-TT beyond compression. Overall, TerraCodec establishes a foundation for exploring the benefits of temporal modeling in EO compression and for advancing multispectral learned compression.
@@ Acknowledgments
Appendix 0.A Limitations
While TerraCodec is pretrained on a large and diverse global dataset, the current models are limited to Sentinel-2 L2A imagery. The underlying architectures and training pipeline are sensor-agnostic, but transferring to other sensors requires finetuning or retraining. From a compression perspective, this reflects a common trade-off: sensor-specific codecs typically achieve higher compression efficiency by modeling the statistics and dynamic range of a particular instrument, whereas sensor-agnostic models offer broader applicability. We see cross-sensor and multisensor pretraining as promising future directions, while also recognizing the practical value of sensor-specific codecs in operational EO pipelines. Our implementation, which is open-sourced, prioritizes research clarity over optimized inference speed; for instance, efficiency techniques such as KV caching are not yet integrated. Incorporating such improvements could further enhance the models’ practicality for real-world deployment.
Our temporal models operate with a two-frame context and are trained and evaluated on four-frame sequences. Using longer sequences would likely reveal additional temporal gains but would also incur substantially higher computational cost. Our P-frame ablation (Fig. 6) highlights the temporal benefits achievable under longer-sequence evaluation.
Finally, FlexTEC establishes a strong flexible-rate baseline, leveraging a latent inference mechanism in which rate control is achieved by dropping latent tokens and inferring the corresponding missing parts through the temporal latent model at decode time. While FlexTEC performs competitively across many settings, there remains room to further narrow the gap to fixed-rate models, particularly under very high compression ratios. Exploring hybrids with other variable-rate methods is a promising future direction.
Appendix 0.B Technical implementation details
@@
0.B.1 Framework and environment
All models are implemented in PyTorch, using CompressAI [7] for core architectures and entropy bottlenecks, extended to support multispectral inputs, temporal samples, and model-belief analyses. The TEC-TT implementations are based on the Neural compression repository [42]. Experiments are run on NVIDIA A100 GPUs with mixed precision training.
0.B.2 Image models (TEC-FP, TEC-ELIC)
TerraCodec-FP follows Factorized Prior [5], with $g_{a}$ and $g_{s}$ implemented as four strided $5{\times}5$ convolutions combined with GDN/IGDN nonlinearities.
TerraCodec-ELIC builds on the uneven channel-group entropy model [26], using the Chandelier reimplementation [13]. Latents are divided into groups [16, 16, 32, 64, M-128] and coded sequentially using SCCTX (space–channel context). Relative to ELIC [26], we omit the preview head for efficiency. Despite the stronger context model, training time remains comparable to TEC-FP since the channel-group autoregression is parallelizable across spatial locations and implemented with efficient masked convolutions.
@@ 0.B.2 Image models (TEC-FP, TEC-ELIC)
Temporal sampling. For image models, we treat the pretraining data as an image dataset by sampling individual timesteps from the SSL4EO-S12 time series. During training, a single temporal index is randomly drawn for every sample in each epoch, such that one epoch covers one quarter of the temporal data. Over the 100 training epochs, this amounts to about 25 full passes through the complete dataset.
0.B.3 Temporal model (TEC-TT)
We provide additional details on the TEC-TT architecture, including its tokenization strategy and temporal transformer design, following VCT [39].
The image encoder–decoder follows the TEC-ELIC backbone, composed of residual bottleneck and attention blocks with a total downsampling factor of $\times 16$. For our $256{\times}256$ input crops, this yields a $16{\times}16$ latent grid with M=192M{=}192 channels.
Tokenization. We tokenize the latent image representations using the scheme introduced in VCT. The latent grid $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{H_{\ell}\times W_{\ell}\times d_{\text{lat}}}$ with $H_{\ell}=W_{\ell}=16$ is divided into spatial blocks. The current frame is split into non-overlapping $4{\times}4$ blocks, each flattened into a sequence of T=16T{=}16 tokens $\{\hat{\mathbf{y}}_{i,b,t}\}_{t=1}^{16}$. The two past frames are partitioned into overlapping $8{\times}8$ context blocks using reflect-padding so their grids align with the current frame, producing T=64T{=}64 tokens per block. Concretely, for block index $b\in\{1,\dots,B\}$:
@@ 0.B.3 Temporal model (TEC-TT)
Temporal transformer stack. The temporal model follows the VCT design and consists of two separate encoders $E_{\text{sep}}$ (one per past frame), a joint encoder $E_{\text{joint}}$ that fuses both contexts, and a masked decoder that autoregressively models the current block tokens conditioned on the fused context. We adopt the standard VCT specifications for the number of layers, heads, and embedding size in each transformer (see Table 4).
@@
Training. Training follows the standard uniform-noise quantization surrogate [52, 41], while inference applies hard quantization and arithmetic coding. We train six TEC-TT variants at different rate–distortion trade-offs, controlled by $\lambda\in\{0.4,5.0,20.0,100.0,300.0,700.0\}$, where smaller values enforce higher compression and larger values prioritize reconstruction quality.
Appendix 0.C Latent Repacking and FlexTEC